Chapter 3, Problem 007 A particle charge of 1.4 yc is at the center of a Gaussien cube 64 cm on edge. What is the net electric flux through the surface? Number Units the tolerance is +/-2%
Answer #1

Solution,

Flux, phi = q/e0 = (1.4 x 10^-6)/(8.85 x 10^-12) = 158192.09 m^2/C

Comment in case any doubt please rate my answer ... ..